5 A pile driver works by lifting a large mass and dropping it to the ground. A driver used to put pilings in the ground for a tall building has a mass of 6150 kg and is raised to a height of 2.5 m in 0.48 s. (a) How much work is done by the engine each time the weight is lifted? (b) What horsepower engine must be used to run the pile driver?
Answer #1

a.)W = mgh = 6150 kg x 9.81m/s2 x 2.5m = 150828.75 J = 150.82875 KJ

b.)P = W/t = 150.82875 Kj/0.48 s = 314.2265 KW

but 1 Kw = 1.34102 hp

so, P = 314.2265x1.34102 hp = 421.384 hp